OpenStudy (anonymous):

radius= 15cm use 3.14 for pie

OpenStudy (anonymous):



pooja195 (pooja195):

What are we finding? Whats the question ? :)

OpenStudy (anonymous):

mmm \(\pi\)

OpenStudy (anonymous):

were tryng to find the area

pooja195 (pooja195):

\[\huge~\rm~\bf~A=\pi~r^2\]

pooja195 (pooja195):

\[\huge~\rm~\bf~A=(3.14)(~15^2)\]

pooja195 (pooja195):

Solve for A

OpenStudy (anonymous):

94.2

pooja195 (pooja195):

hmm can you show me what you did?

OpenStudy (anonymous):

3.14 x15x2

pooja195 (pooja195):

15^2=15 x 15

OpenStudy (anonymous):

lol i forgot im sorry 706.5

pooja195 (pooja195):

706.86 works too :P well done =)

OpenStudy (anonymous):

thanx